Solve for x, y, z, and t in the matrix equation below.  3x y − x t + 1 2 z t − z  =  3 1 7 2 3  .


1
Expert's answer
2021-05-13T01:15:18-0400

To solve for x, y , z, t


"\begin{bmatrix}
 3x & y-x \\
 t+\frac{1}{2}z & t-z
\end{bmatrix}=\begin{bmatrix}
 3 & 1 \\
 \frac{7}{2} & 3
\end{bmatrix}"


On comparing:

"\Rightarrow 3x=3\\\Rightarrow\boxed{x=1}"


"\Rightarrow y-x=1\\\Rightarrow y-1=1\\\Rightarrow \boxed{y=2}"


and

"t+\frac{1}{2}z=\frac{7}{2}\ \ ......(i)\\\ \\t-z=3\ \ .........(ii)\\Subtracting\ equations \ (ii)\ from\ (i)\\We\ get\\\boxed{z=\frac{1}{3}}\\and\ \\t-z=3\\t-\frac{1}{3}=3\\\boxed{t=\dfrac{10}{3}}"